CanPipe Industries purchased a metal stamping machine for CAD 147,000 on March 1, 2019. It is

expected to have useful life of 10 years, salvage value of CAD 27,000, production of 250,000

units, and working hours of 30,000. During 2019 CanPipe used stamping machine for 2450 hours

to stamp 23,450 units of pipeline. From the information given, compute the book depreciation

for 2019 under each of the following methods and comment on your answer:

c. Straight-line method

d. Units of production method

e. Working hours

f. Sum-of-years-digits

g. Declining balance

h. Double declining balance


1
Expert's answer
2021-04-13T10:42:00-0400

c Straight-line method

annual depreciation = (purchase price - salvage value) / useful life

"=\frac{147000-27000}{10}=12000"


d. Units of production method

={(purchase price - salvage value)/estimated total units to be produced }*actual units produced

"=\frac{147000-27000}{250000}\times23450"

"=0.48\times 23450=11256"


e. Working hours

=(purchase price - salvage value)*estimated life of the machine in hours

"=\frac{147000-27000}{30000}=4"


f. Sum-of-years-digits

fractional part of the sum of all years *depreciate cost

depreciate cost"=147000-27000=120000"

sum of all years"=1+2+3+4+5+6+7+8+9+10=55"

10th year"=\frac{1}{55}\times100=1.82" %

"=\frac{1.82}{100}\times120000=2184"


g. Declining balance

basic depreciation rate x book value

depreciation rate"=\frac{100}{10}=10" %

"=\\frac{10}{100}\u200b\u00d7147000=14700"


h. Double declining balance

2*basic depreciation rate x book value

depreciation rate"=\frac{100}{10}=10" %

"=2\times\frac{10}{100}\times 147000=29400"